# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 3 Smoothed Complexity Upper Bound
Definition 25 (Facet and edge event).
@@
Definition 27 (Clockwise order of edges and vertices).
For any given two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ we denote an arbitrary but fixed rotation as “clockwise”. For the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ of our interest, let $p_{1},\dots,p_{k}$ denote its vertices in clockwise order and write $p_{k+1}=p_{1},p_{k+2}=p_{2}$. Then for any edge $e=[p_{i-1},p_{i}]$, we call $p_{i}$ its second vertex in clockwise order and we call $p_{i+1}$ the next vertex after $e$ in clockwise order. The edge $[p_{i},p_{i+1}]$ is the next edge after $e$ in clockwise order.
@@
Theorem 28 (Smoothed complexity upper bound for continuous perturbations).
@@
where we write $[p_{1},p_{2}]=F_{I}\cap W$ and $p_{3}\in\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ as the next vertex after $F_{I}\cap W$ in clockwise order. Then we have
@@
Definition 29 (Exterior angle and length of the next edge).
Given any $I\in\binom{[n]}{d}$, we define two random variables $\theta_{I},\ell_{I^{+}}\geq 0$. If $E_{I}$ happens, write $v\in F_{I}\cap W$ for the second endpoint of $F_{I}\cap W$ in clockwise order. Let $\theta_{I}$ to be the (two-dimensional) exterior angle of $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ at $v$; If $E_{I}$ doesn’t happen then let $\theta_{I}=0$.
Let $\ell_{I^{+}}$ denote the following random variable: If $E_{I}$ happens, then $\ell_{I^{+}}$ equals the length of the next edge after $F_{I}\cap W$ in clockwise order, i.e., the other edge of $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ containing $v$. If $E_{I}$ doesn’t happen then let $\ell_{I^{+}}=0$.
Proof of Theorem 28.
Since we have non-degeneracy with probability $1$, by 23 and linearity of expectation we find
@@
In this case, $\mathbb{E}[\ell_{I^{+}}\mid E_{I}]\geq\frac{t}{20}$, therefore we obtain from $\mathbb{E}[\ell_{I^{+}}]=\mathbb{E}[\ell_{I^{+}}\mid E_{I}]\Pr[E_{I}]$ that
@@
Case 3: $\Pr[E_{I}]>10\binom{n}{d}^{-1}$ and $\Pr[\ell_{I^{+}}\leq t\mid E_{I}]\geq\frac{19}{20}$. Conditional on $E_{I}$, without loss of generality we write $[p_{1},p_{2}]=F_{I}\cap W$ and let $p_{3}$ to denote the next vertex after $F_{I}\cap W$ in clockwise direction. From the theorem’s assumption we have $\Pr[\operatorname{dist}(\operatorname{affhull}(p_{1},p_{2}),p_{3})\geq\gamma\mid E_{I}]\geq\frac{1}{10}$. Then from the union bound,
@@
Since $\theta_{I}\geq 0$ and
@@
we have $\mathbb{E}[\theta_{I}\mid E_{I}]\geq\frac{1}{20}\cdot\frac{\gamma}{t}$, and therefore we can upper bound $\Pr[E_{I}]$ by
@@
Readers are referred to Figure 1 for more illustration of the proof.
@@
To upper bound the second term of (3), we notice that $\sum_{I\in\binom{[n]}{d}}\ell_{I^{+}}$ exactly equals the perimeter of $\operatorname{conv}(a_{1},\dots,a_{n})\cap W$. Since the shadow polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ is contained in the two-dimensional disk of radius $\max_{i\in[n]}\|\pi_{W}(a_{i})\|$, by the monotonicity of surface area for convex sets we have
@@
To upper bound the third term of (3), we notice that the sum of exterior angles for any polygon always equals $2\pi$. Thus
@@
Finally, we combine (3 - 14) and minimize over all $t\geq 0$:
@@
where in the final step, we set $t=\sqrt{\gamma\mathbb{E}[\max_{i\in[n]}\|\pi_{W}(a_{i})\|]}$. ∎
In the subsequent sections, we will show a lower bound for the edge-to-vertex distance $\gamma$ assuming the independently distributed vectors $a_{1},\cdots,a_{n}$ follow Laplace-Gaussian distributions. This allows us to directly apply Theorem 28 to derive an upper bound on the expected number of edges of $\operatorname{conv}(a_{1},\cdots,a_{n})\cap W$. Furthermore, by using Lemma 17, we can further reduce our upper bound to the case when $a_{1},\dots,a_{n}$ are Gaussian distributed vectors.
4 Upper Bound for Two Dimension
In this section, we establish the smoothed complexity upper bound for $d=2$. For this scenario, the shadow plane $W$ encompasses the entire two-dimensional Euclidean space, and $P\cap W$ is identical to $P=\operatorname{conv}(a_{1},\cdots,a_{n})$. From Theorem 28 and Lemma 17, it remains to lower bound the distance from the affine hull of an edge to its neighboring vertex in clockwise order (denoted by $\gamma$ in Theorem 28), where the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})$ is under Laplace-Gaussian perturbation. We will demonstrate a slightly stronger result: a lower bound for the distance between the affine hull of an edge to all of the remaining $(n-2)$ vertices.
Lemma 30 (Edge-to-vertex distance in Two Dimension).
Let $a_{1},\ldots,a_{n}\in\mathbb{R}^{2}$ be $n$ independent $L$-log-Lipschitz random variables. Then for any $I\in\binom{[n]}{2}$, conditional on $E_{I}$ happening, the outer unit normal $y\in W$ of the edge $\operatorname{conv}(a_{i}:i\in I)$ satisfies
@@ 4 Upper Bound for Two Dimension
Together with Theorem 28, Lemma 17, and the Laplace-Gaussian tail bound Lemma 16, we find the upper bound for two-dimensional polygons under Gaussian perturbation:
Theorem 31 (Two-Dimensional Upper Bound).
@@
Proof.
For each $i\in[n]$, let $\hat{a}_{i}$ be independently sampled form the 2-dimensional Laplace-Gaussian distribution $LG_{2}(\bar{a}_{i},\sigma,4\sigma\sqrt{2\log n})$. It follows from Lemma 16 that $\hat{a}_{i}$ is $(4\sigma^{-1}\sqrt{2\log n})$-log-Lipschitz and $\mathbb{E}[\max_{i\in[n]}\|\hat{a}_{i}\|]\leq 1+4\sigma\sqrt{2\log n}$. We use Lemma 30 by setting $L=4\sigma^{-1}\sqrt{2\log n}$, and Theorem 28 by setting $\gamma=\frac{1}{L}=\frac{\sigma}{4\sqrt{2\log n}}$, to find
$\mathbb{E}\left[\operatorname{edges}\left(\operatorname{conv}(\hat{a}_{1},\ldots,\hat{a}_{n})\right)\right]$ $\leq O\left(\frac{\sqrt[4]{\log n}}{\sqrt{\sigma}}+\sqrt{\log n}\right).$
Finally, from Lemma 17, we conclude that $\mathbb{E}\left[\operatorname{edges}\left(\operatorname{conv}(a_{1},\ldots,a_{n})\right)\right]\leq 1+O(\frac{\sqrt[4]{\log n}}{\sqrt{\sigma}}+\sqrt{\log n}).$
Proof of Lemma 30.
@@
Note that $E_{I}$ is now equivalent to either having $z^{\top}a_{j}<t$ for all $j\notin I$ or having $z^{\top}a_{j}>t$ for all $j\notin I$. Write $E_{I}^{+}$ for the former case and $E_{I}^{-}$ for the latter. The vector $z$ is always defined, assuming non-degeneracy, and is equal to the outer normal unit vector $y$ if $E_{I}^{+}$ and equal to $-y$ if $E_{I}^{-}$.
Using Fubini’s theorem, we condition on the values of $a_{j},j\notin I$ and $z$ using Lemma 19. Let $\mu:\mathbb{R}\to\mathbb{R}_{\geq 0}$ denote the induced density of $t=y^{\top}a_{i}=y^{\top}a_{i^{\prime}}$. Then from Lemma 20, $\mu$ is $(2L)$-log-Lipschitz.
@@
$\Pr[(t-\max_{j\notin I}z^{\top}a_{j}\geq\frac{1}{L})\wedge E_{I}^{+}\mid a_{j},j\notin I,z]$
$=$ $\int_{\max_{j\notin I}z^{\top}a_{j}+1/L}^{\infty}\mu(x)\mathrm{d}x$
$=$ $\int_{\max_{j\notin I}z^{\top}a_{j}}^{\infty}\mu(z+1/L)\mathrm{d}x$
$\geq$ $\int_{\max_{j\notin I}z^{\top}a_{j}}^{\infty}e^{-2}\mu(z)\mathrm{d}z$ (By $(2L)$-log-Lipschitzness of $\mu$)
@@
Now observe that, for $i\in I$,
$\Pr[y^{\top}a_{i}-\max_{j\notin I}y^{\top}a_{j}\geq\frac{1}{L}\wedge E_{I}\mid a_{j},j\notin I]$
@@ 5 Multi-Dimensional Upper Bound
In this section, we establish the upper bound for the higher-dimensional case (i.e. $d\geq 3$):
Theorem 32 (Multi-dimensional Upper Bound).
Given any $d>2,n\geq d$, and $\sigma\leq\frac{1}{8d\sqrt{\log n}}$. Let $\bar{a}_{1},\ldots,\bar{a}_{n}$ be $n$ vectors with $\max_{i\in[n]}\|\bar{a}_{i}\|\leq 1$. For each $i\in[n]$, let $a_{i}$ be independently distributed as $\mathcal{N}_{d}(\bar{a}_{i},\sigma^{2}I)$. Then
$\mathbb{E}[\operatorname{edges}(\operatorname{conv}(a_{1},\dots,a_{n})\cap W)]=O\left(\sigma^{-3/2}d^{13/4}\log^{7/4}n\right).$ (15)
Similar to the two-dimensional case (see Section 4), the main technical ingredient of Theorem 32 is a lower-bound of the edge-to-vertex distance (the quantity $\gamma$ in Theorem 28) on the shadow polygon:
Lemma 33 (Edge-to-vertex distance of shadow polygon in multi-dimension).
For any $d\geq 3$, let $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$ be independent $L$-log-Lipschitz random variables. For any $I\in\binom{[n]}{d}$ that satisfies $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1},$ (where $E_{I}$ is defined in Definition 25), we have
@@
Theorem 32 then immediately follows from Lemma 33, Theorem 28, and Lemma 17:
Proof of Theorem 32.
For each $i\in[n]$, let $\hat{a}_{i}$ be independently sampled from the Laplace-Gaussian distribution $LG_{d}(\bar{a}_{i},\sigma,4\sigma\sqrt{d\log n})$. From Lemma 16, we know that
@@
    $\mathbb{E}[\max_{i\in[n]}\|\pi_{W}(\hat{a}_{i})\|]\leq 1+4\sigma\sqrt{2\log n}\leq 1.5$.
From Lemma 33, we get that for any $p\in F_{I}\cap W$, if $p^{\prime}$ is the next vertex after the edge $F_{I}\cap W$ in clockwise order, then
$\Pr[y_{I}^{\top}p\geq y_{I}^{\top}p^{\prime}+\Omega(\frac{1}{L^{3}d^{5}\log^{2}n})\mid E_{I}]\geq 0.1.$
Here $y_{I}\in W$ is the outer unit normal vector of the polygon $\operatorname{conv}(\hat{a}_{1},\ldots,\hat{a}_{n})\cap W$ on the edge $F_{I}\cap W$. Then we can use Theorem 28 by setting $L=4\sigma^{-1}\sqrt{d\log n}$, $\gamma=\Omega(\frac{1}{L^{3}d^{5}\log^{2}n})$ and $\mathbb{E}[\max_{i\in[n]}\|\pi_{W}(a_{i})\|]=1.5$, to find
$\mathbb{E}\left[\operatorname{edges}\left(\operatorname{conv}(\hat{a}_{1},\ldots,\hat{a}_{n})\right)\right]\leq$ $10+O(\sqrt{\sigma^{-3}d^{13/2}\log^{7/2}n}).$
Finally, from Lemma 17, we conclude that
$\mathbb{E}\left[\operatorname{edges}\left(\operatorname{conv}(a_{1},\ldots,a_{n})\cap W\right)\right]$ $\leq 11+O(\sqrt{\sigma^{-3}d^{13/2}\log^{7/2}n})$
$=O\left(\sigma^{-3/2}d^{13/4}\log^{7/4}n\right).$
The rest of this section will be structured as follows. In section 5.1 we define some basic notations that will be used in the proof. In section 5.2 we provide a deterministic argument to establish sufficient criteria for the conclusion of Lemma 33 to hold. In section 5.3 and section 5.4, we prove that these conditions hold with good probability conditional on $E_{I}$. The proof of Lemma 33 is then finished in section 5.5.

### v3

@@ 3 Smoothed Complexity Upper Bound
Definition 3.1 (Facet and edge event).
@@
Definition 3.3 (Clockwise order of edges and vertices).
For any given two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ we denote an arbitrary but fixed rotation as “clockwise”. For the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ of our interest, let $p_{1},\dots,p_{k}$ denote its vertices in clockwise order and write $p_{k+1}=p_{1},p_{k+2}=p_{2}$. Then for any edge $e=[p_{i-1},p_{i}]$, we call $p_{i}$ its second vertex in clockwise order and we call $p_{i+1}$ the next vertex after $e$ in clockwise order. The edge $p_{i},p_{i+1}]$ is the next edge after $e$ in clockwise order.
@@
Theorem 3.4 (Smoothed complexity upper bound for continuous perturbations).
@@
where we write $p_{1},p_{2}]=F_{I}\cap W$ and $p_{3}\in\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ as the next vertex after $F_{I}\cap W$ in clockwise order. Then we have
@@
Definition 3.5 (Exterior angle and length of the next edge).
Given any $I\in\binom{[n]}{d}$, we define two random variables $\theta_{I},\ell_{I^{+}}\geq 0$. If $E_{I}$ happens, write $p_{1},p_{2}\in F_{I}\cap W$ for the first and the second endpoint of $F_{I}\cap W$ in clockwise order. Let $\theta_{I}$ to be the (two-dimensional) exterior angle of $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ at $p_{2}$; If $E_{I}$ doesn’t happen then let $\theta_{I}=0$.
Let $\ell_{I^{+}}$ denote the following random variable: If $E_{I}$ happens, then $\ell_{I^{+}}$ equals to the length of the next edge after $F_{I}\cap W$ in clockwise order, i.e., the other edge of $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ containing $p_{2}$. If $E_{I}$ doesn’t happen then let $\ell_{I^{+}}=0$.
Proof 3.6 (Proof of Theorem 3.4).
Since we have non-degeneracy with probability $1$, by Fact 4 and linearity of expectation we find
@@
In this case, by Markov’s inequality $\mathbb{E}[\ell_{I^{+}}\mid E_{I}]\geq\frac{t}{20}$, therefore we obtain from $\mathbb{E}[\ell_{I^{+}}]=\mathbb{E}[\ell_{I^{+}}\mid E_{I}]\Pr[E_{I}]$ and $E[\ell_{I^{+}}|E_{I}^{C}]=0$ that
@@
Case 3: $\Pr[E_{I}]>10\binom{n}{d}^{-1}$ and $\Pr[\ell_{I^{+}}\leq t\mid E_{I}]\geq\frac{19}{20}$. Readers are referred to Figure 1 for more illustration of the proof.
Conditional on $E_{I}$, without loss of generality we write $p_{1},p_{2}]=F_{I}\cap W$ and let $p_{3}$ denote the next vertex after $F_{I}\cap W$ in clockwise direction. From the theorem’s assumption we have $\Pr[\operatorname{dist}(\operatorname{affhull}(p_{1},p_{2}),p_{3})\geq\gamma\mid E_{I}]\geq\frac{1}{10}$. Then from the union bound,
@@
Referring back to Figure 1, we know that $\theta_{I}\geq 0$ and thus
@@
we have $\mathbb{E}[\theta_{I}\mid E_{I}]\geq\frac{1}{20}\cdot\frac{\gamma}{t}$. Combining this with the fact that $\theta_{I}=0$ if $E_{I}$ does not hold, we know that $\mathbb{E}[\theta_{I}]=\mathbb{E}[\theta_{I}\mid E_{I}]\Pr[E_{I}]$ and we can upper bound
@@
To upper bound the second term of (3.6), we notice that $\sum_{I\in\binom{[n]}{d}}\ell_{I^{+}}$ exactly equals the perimeter of $\operatorname{conv}(a_{1},\dots,a_{n})\cap W$. Since the shadow polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ is contained in the two-dimensional disk of radius $\max_{i\in[n]}\|\pi_{W}(a_{i})\|$, by the monotonicity of surface area for convex sets we have
@@
To upper bound the third term of (3.6), we notice that the sum of exterior angles for any polygon always equals $2\pi$. Thus
@@
Finally, we combine (3.6 - 14) and minimize over all $t>0$:
@@
where in the final step, we set $t=\sqrt{\gamma\mathbb{E}[\max_{i\in[n]}\|\pi_{W}(a_{i})\|]}$.
In the subsequent sections, we will show a lower bound for the edge-to-vertex distance $\gamma$ assuming the independently distributed vectors $a_{1},\cdots,a_{n}$ follow Laplace-Gaussian distributions. This allows us to directly apply Theorem 3.4 to derive an upper bound on the expected number of edges of $\operatorname{conv}(a_{1},\cdots,a_{n})\cap W$. Furthermore, by using Lemma 2.11, we can further reduce our upper bound to the case when $a_{1},\dots,a_{n}$ are Gaussian distributed vectors.
4 Upper Bound in Two Dimension
In this section, we establish the smoothed complexity upper bound for $d=2$. For this scenario, the shadow plane $W$ encompasses the entire two-dimensional Euclidean space, and $P\cap W$ is identical to $P=\operatorname{conv}(a_{1},\cdots,a_{n})$. From Theorem 3.4 and Lemma 2.11, it remains to lower bound the distance from the affine hull of an edge to its neighboring vertex in clockwise order (denoted by $\gamma$ in Theorem 3.4), where the vertices of the polygon $a_{1},\cdots,a_{n}$ is sampled from a Laplace-Gaussian distribution with the center of $\bar{a}_{1},\cdots,\bar{a}_{n}$. We will demonstrate a slightly stronger result: a lower bound for the distance between the affine hull of an edge to all of the remaining (n2)(n-2) vertices.
Lemma 4.1 (Edge-to-vertex distance in Two Dimension).
Let $a_{1},\ldots,a_{n}\in\mathbb{R}^{2}$ be $n$ independent $L$-log-Lipschitz random variables. Then, for any $I\in\binom{[n]}{2}$, the outer unit normal $y\in W$ of the edge $\operatorname{conv}(a_{i}:i\in I)$ satisfies
@@ 4 Upper Bound in Two Dimension
By applying Theorem 3.4, Lemma 2.11, and the Laplace-Gaussian tail bound of Lemma 2.10, we find the following upper bound for two-dimensional polygons under Gaussian perturbation:
Theorem 4.2 (Two-Dimensional Upper Bound).
@@
Proof 4.3.
For each $i\in[n]$, let $\hat{a}_{i}$ be independently sampled form the 2-dimensional Laplace-Gaussian distribution $LG_{2}(\bar{a}_{i},\sigma,4\sigma\sqrt{2\log n})$. It follows from Lemma 2.10 that $\hat{a}_{i}$ is $(4\sigma^{-1}\sqrt{2\log n})$-log-Lipschitz and $\mathbb{E}[\max_{i\in[n]}\|\hat{a}_{i}\|]\leq 1+4\sigma\sqrt{2\log n}$. We use Lemma 4.1, setting $L=4\sigma^{-1}\sqrt{2\log n}$, and Theorem 3.4, setting $\gamma=\frac{1}{L}=\frac{\sigma}{4\sqrt{2\log n}}$, to find
$\mathbb{E}\left[\operatorname{edges}\left(\operatorname{conv}(\hat{a}_{1},\ldots,\hat{a}_{n})\right)\right]$ $\leq O\left(\sqrt{\frac{\sqrt{\log n}}{\sigma}+\log n}\right)\leq O\left(\frac{\sqrt[4]{\log n}}{\sqrt{\sigma}}+\sqrt{\log n}\right).$
Finally, from Lemma 2.11, we conclude that $\mathbb{E}\left[\operatorname{edges}\left(\operatorname{conv}(a_{1},\ldots,a_{n})\right)\right]\leq 1+O(\frac{\sqrt[4]{\log n}}{\sqrt{\sigma}}+\sqrt{\log n}).$
Proof 4.4 (Proof of Lemma 4.1).
@@
Note that $E_{I}$ is now equivalent to either having $z^{\top}a_{j}<t$ for all $j\notin I$ or having $z^{\top}a_{j}>t$ for all $j\notin I$. Write $E_{I}^{+}$ for the former case and $E_{I}^{-}$ for the latter. The vector $z$ is always defined, assuming non-degeneracy, and is equal to the outer normal unit vector $y$ conditional on $E_{I}^{+}$ and equal to y-y conditional on $E_{I}^{-}$.
We want to apply the principle of deferred decisions to fix the values of $a_{j}$ for each $j\notin I$ and $z$, and now only allow $a_{i}$ and $a_{i^{\prime}}$ to vary while fixing the slope of the line between $a_{i}$ and $a_{i^{\prime}}$. Let $\mu:\mathbb{R}\to\mathbb{R}_{\geq 0}$ denote the induced probability density function of $t=y^{\top}a_{i}=y^{\top}a_{i^{\prime}}$. Lemma 2.14 tells us that $\mu$ is $(2L)$-log-Lipschitz.
@@
$\Pr[(t-\max_{j\notin I}z^{\top}a_{j}\geq\frac{1}{L})\wedge E_{I}^{+}]$
$=$ $\int_{\max_{j\notin I}z^{\top}a_{j}+1/L}^{\infty}\mu(t)\mathrm{d}t$
$=$ $\int_{\max_{j\notin I}z^{\top}a_{j}}^{\infty}\mu(s+1/L)\mathrm{d}s$
$\geq$ $\int_{\max_{j\notin I}z^{\top}a_{j}}^{\infty}e^{-2}\mu(s)\mathrm{d}s$ (By $(2L)$-log-Lipschitzness of $\mu$)
@@
Now observe that, for any $i\in I$, we have
$\Pr[y^{\top}a_{i}-\max_{j\notin I}y^{\top}a_{j}\geq\frac{1}{L}\wedge E_{I}]$
@@ 5 Multi-Dimensional Upper Bound
In this section, we establish the upper bound for the higher-dimensional case (i.e., $d\geq 3$):
Theorem 5.1 (Multi-dimensional Upper Bound).
Let $d>2,n\geq d$, and $\sigma\leq\frac{1}{16d\sqrt{\log n}}$. Let $\bar{a}_{1},\ldots,\bar{a}_{n}$ be $n$ vectors with $\max_{i\in[n]}\|\bar{a}_{i}\|\leq 1$. For each $i\in[n]$, let $a_{i}$ be independently distributed as $\mathcal{N}_{d}(\bar{a}_{i},\sigma^{2}I)$. Then
$\mathbb{E}[\operatorname{edges}(\operatorname{conv}(a_{1},\dots,a_{n})\cap W)]=O\left(\sigma^{-3/2}d^{13/4}\log^{5/4}n\right).$ (15)
Similar to the two-dimensional case (see Section 4), the main technical ingredient of Theorem 5.1 is a lower-bound of the edge-to-vertex distance (the quantity $\gamma$ in Theorem 3.4) on the shadow polygon:
Lemma 5.2 (Edge-to-vertex distance of shadow polygon in multi-dimension).
For any $d\geq 3$, let $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$ be independent $L$-log-Lipschitz random variables. For any $I\in\binom{[n]}{d}$ that satisfies $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1},$ (where $E_{I}$ is defined in Definition 3.1), we have
@@
Theorem 5.1 then immediately follows from Lemma 5.2, Theorem 3.4, and Lemma 2.11:
Proof 5.3 (Proof of Theorem 5.1).
For each $i\in[n]$, let $\hat{a}_{i}$ be independently sampled from the Laplace-Gaussian distribution $LG_{d}(\bar{a}_{i},\sigma,4\sigma\sqrt{d\log n})$. From Lemma 2.10, we know that
@@
    $\mathbb{E}[\max_{i\in[n]}\|\pi_{W}(\hat{a}_{i})\|]\leq 1+8\sigma\sqrt{2\log n}\leq 1.5$.
From Lemma 5.2, we get that for any $p\in F_{I}\cap W$, if $p^{\prime}$ is the next vertex after the edge $F_{I}\cap W$ in clockwise order, then
$\Pr[y_{I}^{\top}p\geq y_{I}^{\top}p^{\prime}+\Omega(\frac{1}{L^{3}d^{5}\log n})\mid E_{I}]\geq 0.1.$
Here $y_{I}\in W$ is the outer unit normal vector of the polygon $\operatorname{conv}(\hat{a}_{1},\ldots,\hat{a}_{n})\cap W$ on the edge $F_{I}\cap W$. Then we can use Theorem 3.4 by setting $L=4\sigma^{-1}\sqrt{d\log n}$, $\gamma=\Omega(\frac{1}{L^{3}d^{5}\log n})$ and $\mathbb{E}[\max_{i\in[n]}\|\pi_{W}(a_{i})\|]=1.5$, to find
$\mathbb{E}\left[|\operatorname{edges}\left(\operatorname{conv}(\hat{a}_{1},\ldots,\hat{a}_{n})\right)|\right]$ $\leq 10+O(\sqrt{\frac{E[\max_{i\in[n]}\|\pi_{W}(a_{i})\|]}{\gamma}})$
$\leq 10+O(\sqrt{\frac{1.5}{\frac{1}{L^{3}d^{5}\log n}}})$
$\leq 10+O(\sqrt{L^{3}d^{5}\log n})$
$\leq 10+O(\sqrt{\sigma^{-3}d^{13/2}\log^{5/2}n}).$
Finally, from Lemma 2.11, we conclude that
$\mathbb{E}\left[|\operatorname{edges}\left(\operatorname{conv}(a_{1},\ldots,a_{n})\cap W\right)|\right]$ $\leq 11+O(\sqrt{\sigma^{-3}d^{13/2}\log^{5/2}n})$
$=O\left(\sigma^{-3/2}d^{13/4}\log^{5/4}n\right).$
The rest of this section is dedicated to the proof of Lemma 5.2 and will be structured as follows. In Section 5.1 we define some basic notation that will be used in the proof. In Section 5.2 we establish two sufficient criteria for the conclusion of Lemma 5.2 to hold. In Section 5.3 and Section 5.4, we prove that these conditions hold with good probability conditional on $E_{I}$. Section 5.5 to Section 5.7 include the proof of the auxiliary lemmas. Finally, we finish the proof of Lemma 5.2 in Section 5.8.